OpenStudy (anonymous):

f(x)=3x^3lnx find f'(e^3)?

OpenStudy (lgbasallote):

\[f(x) = 3x^3 \ln x\] do you know the value of f' (x)?

OpenStudy (anonymous):

its 3x^2(3log(x)+1)

OpenStudy (lgbasallote):

you used product rule here right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

hmm okay..i suggest using ln instead of log though..especially for this problem so f'(e^3) =\(3(e^3)^2(3\ln (e^3) + 1)\)

OpenStudy (lgbasallote):

im guessing your problem was \(\log (e^3)?\) you didnt use ln?